What is the x-intercept in this equation: 3x+2y-10=0
Found 2 solutions by ewatrrr, rfer:
Answer by ewatrrr(24785)


Hi,
the x-intercept for equation:
3x+2y-10=0 |value of x when y=0
3x = 10
x = 10/3 Pt(10/3,0) the x-intercept

Answer by rfer(16322)

3x+2y=10
3x+2*0=10
x=10/3
(10/3,0)